OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 2x − 3y − z if x = −3, y = 2, and z = −2

OpenStudy (anonymous):

@Noemi95 help

OpenStudy (anonymous):

@primeralph help

OpenStudy (anonymous):

@ajprincess help

OpenStudy (anonymous):

plug in known values: 2(-3)-3(2)-(-2)

OpenStudy (raden):

2x − 3y − z = 2(-3) - 3(2) - (-2) = -6 - 6 + 2 = ...

OpenStudy (anonymous):

yup @RadEn 's got it Julian

OpenStudy (anonymous):

is -10 the answers ?

OpenStudy (raden):

yup that's right

OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 3x − 2y if x = 2 and y = −2 @raden

OpenStudy (primeralph):